PREX/CREX Jan 15, 2019 10:00am

From PREX Wiki
Jump to: navigation, search

Back to Main Page >> PREX/CREX_Teleconferences

previous meeting << >> following meeting

Logistic information

 BlueJeans calling instructions:
 Toll-Free Number (U.S.&  Canada):  888-240-2560
 International toll number:         408-740-7256
 Bluejeans CODE:                  628 981 532
 Bluejeans link: https://bluejeans.com/628981532
 Offline room:   CEBAF Center F324-325

Agenda

  1. Design Status: Kent and Ciprian
    • CG: design is mostly complete with the exception of two items
    • CG: the sieve movement mechanism is being designed now and as far as I can tell there are no major concerns on our end that they are not handling (i.e. making sure we know the position of the fully retracted sieves through some physical stops and making sure the mechanism is smooth and not prone to locking).
    • CG: Q1 collimator KP: Seamus used the same z position as before and we should be fine since they are placing the collimator in front of the bar. SR: the output pictures from simulations also matched with what Adam drew.
  2. March 7-10 planning: Don and Ciprian googleDocPlan
    • DJ: assume 4-5 shifts for Moller polarimeter commissioning
    • DJ: first shift is to check the SC coil. Hopefully the alignment is good enough to continue the work over the weekend. If we need to align it is unlikely we will get Survey and Alignment to work Friday evening/night
    • DJ: probably break Moller studies after 24 h to allow other beam studies and come back after 1/2 other shift(s).
    • GU: what is the tolerance for continuing? DJ: that would be up to Yves and Ops.
    • KP: we should plan to either have 2 teams of experts or have at most 2 shifts in a row
  3. Compton update: Adam docDB294
    • AZ: The RRPD signal problem does not seem to be an issue as of this morning. It is not clear why this changed over the last week.
    • DG: plan is still to perform systematic checks to properly understand what is going on on the table
  4. Injector Status: Amali and Caryn docDB295
    • CP: in the next week we will look at the electron beam with chopper scans.
    • CP: we measured the DOCP and got 99.8%. the IHWP was 2deg off and after fixing it we are at 99.9%.
    • DG: you should try to replace the banana cables for the critical spare.
    • CC: we saw a similar issue in the test lab and will talk to Bob to investigate.
  5. Schedule for next mtg: Ciprian Weekly Meeting Discussion Topics
    • Moller update: Eric
    • Integrating beam monitors: Ciprian
    • DAQ: Bob
    • GEMs: Devi
    • Tracking and optics: Chandan and Seamus

Present

In F345

Adam Z, Chandan G, Dave G, Amali P, Sanghwa P, Ciprian G, Eric K, Victoria O, Tao Y, Cameron C, Caryn P,

Online

Devi A, Ryan R, Ye T, Paul K, Dustin M, Guido U, Kent P, Paul S, Seamus R, Jay B, Don J, Jinlong Z, Sakib R,

Excused

KK,